# Ab initio spin Hamiltonians and magnetism of Ce and Yb triangular-lattice compounds

### v1

@@ Abstract
Upon tuning both $J_{\pm\pm}$ and $J_{z\pm}$, $\mathcal{S}(\bm{k})$ develops a peak at the $M$ points, signaling the onset of the stripe phase, as shown in panels $c)$ and $d)$. From the phase diagram in panel (a), we find that along the $J_{z\pm}=0$ line, negative values of $J_{\pm\pm}/J$ favor a stripe-$\parallel$ GS, whereas positive values of $J_{\pm\pm}/J$ stabilize a stripe-$\perp$ state. The coupling $J_{z\pm}$ further extends the stripe-$\perp$ region into negative $J_{\pm\pm}/J$ values, which agrees with previous studies. In certain parameter regimes, the peak in $\mathcal{S}(\bm{k})$ appears in both the $M_{0}$ and $M_{1}$ wave vectors, while in others it shifts to the $M_{2}$ wave vector. This shift does not signify a transition between the two stripe phases, but instead arises from the geometry of the $32$-site cluster (being absent in the $12$-site cluster). As the $32$-site cluster does not have the $C_{3}$ rotation symmetry, the three $M$ points are inequivalent. As a result, the degeneracy among the six stripe configurations is lifted by the lattice geometry. To account for this, in panel (c) of Fig. 3, we display
$\tilde{\mathcal{S}}(M)=[\mathcal{S}(M_{0})+\mathcal{S}(M_{1})+\mathcal{S}(M_{2})]/3$ (4)
averaged over the three $M$ points. Distinguishing between the two stripe types requires examining additional spin-spin correlators, which is left to the SM.
@@
We now set $J_{z\pm}=0$ and study the phase diagram as a function of the NNN coupling $J^{\prime}$ and $J_{\pm\pm}$. Based on the analysis of the quantum numbers of the lowest excited states, we map out the phase diagram identifying four distinct regions, as shown in panel $d)$ of Fig. 3. For $J^{\prime}/J\sim[0.0,0.05]J$, the GS is in an $120^{\circ}$ AFM state, which is further corroborated by the peak in $\mathcal{S}(\bm{k})$ at the $Z$ points (shown in panel $e)$). The $120^{\circ}$ AFM is stabilized up to values of $J_{\pm\pm}\sim 0.1J$. Increasing $J_{\pm\pm}$ promotes the development stripe order, whose precise nature depends on the sign of $J_{\pm\pm}/J$, as shown in panel $d)$. Negative values favor a stripe-$\parallel$ GS, whereas positive values stabilize a stripe-$\perp$. The stripe order is further corroborated by the strong intensity of $\mathcal{S}(\bm{k})$ at the $M$ points, as shown in the panel $f)$. The black and white lines indicating the phase boundaries in Fig. 3 are obtained from the position of level crossings in the excitation spectrum, see SM [56] for details.
In the intermediate regime, $J^{\prime}\sim[0.05,0.16]J$ and $J_{\pm\pm}\sim\pm 0.1J$, the first excited state outside the $\Gamma.\text{A}$ sector transforms according to the $X_{1}.\text{A}$ irrep, as shown in the supplementary material. This parameter range corresponds to the orange region in the phase diagram (panel $d)$). In the context of triangular-lattice antiferromagnets, the appearance of an excitation at momentum $X$ is particularly noteworthy, as it has been interpreted as a signature of a DSL [19, 20]. Specifically, Ref. [19] predicted that the singlet-monopole excitation of the DSL carries momentum $X$ and appears in the low-energy spectrum of the $J_{1}$$J_{2}$ triangular-lattice antiferromagnet in the spin-liquid regime [11]. An alternative explanation for such a low-lying excitation is a 12-site valence-bond solid, for which consistent correlations have also been observed in the spin-liquid regime [11]. In this parameter window, the static spin structure factor exhibits nearly the same intensity at the $Z$ and $M$ points. We further find that this region is stabilized only for small values of $J_{\pm\pm}\sim\pm 0.1J$, while larger values drive the system towards the stripe phase.
@@
Conclusions – We have derived from first principles the crystal-field splitting, Kramer’s GS and intersite exchange interactions within the GS manifold for a set of rare-earth triangular materials, including two systems (KCeO2 and KCeS2) that order magnetically, as well as, two systems (RbCeO2 and YZGO) that have been reported as putative QSLs. The calculated crystal-field splitting and gyromagnetic $g$-factors are in good agreement with experiment. The mean-field Curie-Weiss temperature evaluated from the resulting low-energy Hamiltonian including both the calculated intersite exchange and classical dipole-dipole magnetic interaction agrees well with experiment for all three oxide systems but is severely underestimated for KCeS2.
@@
Solving the resulting low-energy Hamiltonian with ED, we find evidence for the $120^{\circ}$ AFM order in all systems apart from KCeS2, where a stripe-$\perp$ AFM state is instead stabilized, consistent with experimental observations. We find strong evidence for a magnetic $120^{\circ}$ order in the two DSL candidates. From the extended phase diagram of their Hamiltonian, we concluded that a stronger antiferromagnetic NNN coupling is required to realize the DSL phases in comparison to those present in the compounds.
@@
Our calculated intersite exchange interactions include all $f$-$f$ kinetic exchange contributions (in particular, superexchange). However, in rare-earth semiconductors there can be another type of $f$-$f$ intersite exchange generated through 4ff-5dd hopping and the inter-shell Hund’s rule interaction [77]. This could become important in the case of very weak superexchange like in KCeS2, where our calculations apparently reproduce the structure of intersite exchange but not its magnitude. Further studies to evaluate the importance of such exchange mechanisms in triangular rare-earth compounds are worthwhile.

### v2

@@ Abstract
Upon tuning both $J_{\pm\pm}$ and $J_{z\pm}$, $\mathcal{S}(\bm{k})$ develops a peak at the $M$ points, signaling the onset of the stripe phase (see panel $c)$) We find that along the $J_{z\pm}=0$ line, negative values of $J_{\pm\pm}/J$ favor a stripe-$\parallel$, whereas positive values stabilize a stripe-$\perp$ state. The coupling $J_{z\pm}$ further extends the stripe-$\perp$ region into negative $J_{\pm\pm}/J$ values, which agrees with previous studies. In certain parameter regimes, the peak in $\mathcal{S}(\bm{k})$ appears at both $M_{0}$ and $M_{1}$, while in others it shifts to $M_{2}$. This shift does not signal a transition between stripe phases, but originates from the absence of $C_{3}$ symmetry in the $32$-site cluster, which renders the three $M$ points inequivalent (unlike in the $12$-site cluster). Consequently, the lattice geometry lifts the degeneracy among the six stripe configurations. To account for this, in Fig. 3 $c)$ we plot $\mathcal{S}(\bm{k})$ averaged over the three $M$ points,
$\tilde{\mathcal{S}}(M)=\frac{\mathcal{S}(M_{0})+\mathcal{S}(M_{1})+\mathcal{S}(M_{2})}{3}.$ (3)
Distinguishing the two stripe types requires examining additional spin-spin correlators, which we do in the SM. We observe in panel $b)$ that the couplings determined for the YZGO system place it within the $120^{\circ}$ AFM region as expected.
@@
We set $J_{z\pm}=0$ and study the phase diagram as a function of $J^{\prime}$ and $J_{\pm\pm}$, identifying four distinct regions, as shown in Fig. 3 $d)$. For $J^{\prime}\in[0,0.05]J$, the GS is in a $120^{\circ}$ AFM state, confirmed by the peak of $\mathcal{S}(\bm{k})$ at the $Z$ points (panel $e)$). This state persists up to $J_{\pm\pm}\sim 0.1J$. Increasing $J_{\pm\pm}$ induces stripe order, whose orientation depends on the sign of $J_{\pm\pm}$: negative values yield stripe-$\parallel$, while positive values stripe-$\perp$ GS. This is corroborated by the peak of $\mathcal{S}(\bm{k})$ at the $M$ points as shown in panel $f)$.
For $J^{\prime}\sim[0.05,0.16]J$ and $J_{\pm\pm}\sim[-0.1,0.1]J$, the first non-zero momentum excitation belongs to the $X_{1}.\text{A}$ irrep (see SM), corresponding to the orange region in panel $d)$. As discussed, the excitation at momentum $X$ is particularly noteworthy, as it is interpreted as a signature of the singlet-monopole of a DSL [19, 20], predicted to appear in the low-energy spectrum of the $J_{1}$$J_{2}$ triangular-lattice antiferromagnet in the spin-liquid regime [11]. An alternative explanation for such a low-lying excitation is a 12-site valence-bond solid, for which consistent correlations have also been observed in the spin-liquid regime [11]. We find that this region is stabilized only for small values of $J_{\pm\pm}\sim\pm 0.1J$, while larger values drive the system towards the stripe phase. This analysis indicates that the previously reported quantum paramagnetic phases in earlier DMRG studies [54] (identified even in the absence of NNN coupling) might not correspond to a DSL phase. In the SM, we provide additional results of the energy spectra, which reveal that the $X.\text{A}$ states persist at higher energies, whereas the low-lying excitations belong to the tower of states characteristic of the $120^{\circ}$ AFM order. While these results do not definitively exclude other paramagnetic states, the sharp $K$-point peak in the spin structure factor (computed also for an $N=36$ cluster at $J^{\prime}=0$, $J_{z\pm}=0.3J$, and $J_{\pm\pm}=-0.08J$) strongly supports $120^{\circ}$ AFM order. Although a DSL may exhibit similar quasi–long-range signatures, which finite-size ED cannot unambiguously distinguish from true long-range order, our spectral analysis offers little evidence for such a phase at this parameter point.
@@
Conclusions – We have derived from first principles the crystal-field splitting and and intersite exchange interactions within the Kramers GS manifold for a set of rare-earth triangular materials, including two systems (KCeO2 and KCeS2) that order magnetically, as well as two putative QSLs (RbCeO2 and YZGO). The calculated CF splitting and CF GS $g$-factors are in good agreement with experiment. The mean-field Curie-Weiss temperature evaluated from the resulting low-energy Hamiltonian agrees well with experiment for all three oxide systems but is severely underestimated for KCeS2.
@@
Solving the resulting low-energy Hamiltonian with ED, we find evidence for the $120^{\circ}$ order in all systems apart from KCeS2, where a stripe-$\perp$ state is instead stabilized, consistent with experimental observations. We find strong evidence for a magnetic $120^{\circ}$ order in the two DSL candidates. From the extended phase diagram of their Hamiltonian, we concluded that a stronger antiferromagnetic NNN coupling is required to realize the DSL phases in comparison to those present in the compounds.
@@
Our calculated intersite exchange interactions include all $f$-$f$ kinetic exchange contributions (in particular, superexchange). However, in rare-earth semiconductors there can be another type of $f$-$f$ intersite exchange generated through 4ff-5dd hopping and the inter-shell Hund’s rule interaction [76]. This mechanism could become important in the case of very weak superexchange like in KCeS2, where our calculations apparently reproduce the structure of intersite exchange but not its magnitude. Its impact on the NNN exchange $J^{\prime}$ also needs to be evaluated, especially in the light of our present predictions for $J^{\prime}$ to be too small to stablize a DSL phase in YZGO. Further studies to evaluate the importance of such exchange mechanisms in triangular rare-earth compounds are worthwhile.